OpenStudy (anonymous):

Given the polynomial 6x3 + 4x2 − 6x − 4, what is the value of the coefficient 'k' in the factored form? 6x3 + 4x2 − 6x − 4 = 2(x + k)(x − k)(3x + 2) k= ____________ Numerical Answers Expected! Answer for Blank 1:

OpenStudy (anonymous):

\[6 x^3+4x^2-6x-4=2[3x^3+2x^2-3x-2]=2[(3x^3+2x^2)-1(3x+2)]\] \[=2[x^2(3x+2)-1(3x+2)]\] =?

OpenStudy (anonymous):

im sorry i really dont understand how to do any of this..

OpenStudy (amistre64):

the simplest process is to just expand the right side, and compare coefficients

OpenStudy (loser66):

@joannaayala17 follow @amistre64, please

OpenStudy (amistre64):

as a test, lets try this: let x=0 what do we end up with?

OpenStudy (amistre64):

0+0+0-4 = 2(0+k)(0-k)(0+2)

OpenStudy (anonymous):

It's a polynomial of the form \[ax^3+bx^3+c+d\] Consider the "d" term on both the sides Now obviously on the left side this term is -4 on the right side it, it can be obtained be multiplying all constants together, note k is a constant and not a variable \[-4=2 \times k \times (-k) \times2\]

OpenStudy (anonymous):

sorry \[ax^3+bx^2+cx+d\] is what I mean

OpenStudy (anonymous):

If you're having trouble you can always expand all the parenthesis

OpenStudy (anonymous):

and then compare